# Preference Elicitation for Offline Reinforcement Learning

### v1

@@ A.5.2 Analysis of the reward term
Let $\Pi_{offline}$ denote the following set of near-optimal pessimistic policies, under the pessimitic transition model $\hat{T}_{inf}$ and the reward confidence set $\mathcal{R}$:
$\Pi_{offline}=\{\pi\>|\>\pi=\text{argmax}_{\pi\in\Pi}\mathbb{E}_{\tau\sim d^{\pi}_{\hat{T}_{inf}}}\big{[}\tilde{R}(\tau)\big{]}\>\forall\tilde{R}\in\mathcal{R}\}$
Under event $\mathcal{E}_{R}$, we have $\pi^{*}_{offline}\in\Pi_{offline}$.
@@
Recall the definition of $\pi^{*}_{offline}$: $\pi^{*}_{offline}=\text{argmax}_{\pi\in\Pi}V_{\hat{T}_{inf},R}^{\pi}$. Note that there is no need to consider the preference baseline term in $V^{\pi}$ when building $\Pi_{offline}$ since it is independent of the policy. Under event $\mathcal{E}_{R}$, we have $R\in\mathcal{R}$. Thus, $\pi^{*}_{offline}\in\Pi_{offline}$. ∎
@@
$V^{\pi^{*}_{offline}}_{\hat{T}_{inf},R}-V^{\pi^{*}_{offline}}_{\hat{T}_{inf},\hat{R}_{inf}}\leq 2\sqrt{\kappa^{2}c_{R}/N_{p}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}$
@@
$V^{\pi^{*}_{offline}}_{\hat{T}_{inf},R}$ $-V^{\pi^{*}_{offline}}_{\hat{T}_{inf},\hat{R}_{inf}}$
$=(V^{\pi^{*}}_{\hat{T}_{inf},R}-V^{\pi^{*}}_{\hat{T}_{inf},\hat{R}})+(V^{\pi^{*}}_{\hat{T}_{inf},\hat{R}}-V^{\pi^{*}}_{\hat{T}_{inf},\hat{R}_{inf}})$
$=\mathbb{E}_{\tau\sim d^{{\pi}^{*}_{offline}}_{\hat{T}_{inf}}}[R(\tau)]-\mathbb{E}_{\tau\sim d_{pref}}[{R}(\tau)]-\mathbb{E}_{\tau\sim d^{{\pi}^{*}_{offline}}_{\hat{T}_{inf}}}[\hat{R}_{inf}(\tau)]+\mathbb{E}_{\tau\sim d_{pref}}[\hat{R}_{inf}(\tau)]$
$=\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}_{offline}}_{\hat{T}_{inf}},\tau_{2}\sim d_{pref}}[R(\tau_{1})-R(\tau_{2})]-\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}_{offline}}_{\hat{T}_{inf}},\tau_{2}\sim d_{pref}}[\hat{R}_{inf}(\tau_{1})-\hat{R}_{inf}(\tau_{2})]$
$\leq\kappa\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}_{offline}}_{\hat{T}_{inf}},\tau_{2}\sim d_{pref}}[P_{R}(\tau_{1}\succ\tau_{2})-P_{\hat{R}_{inf}}(\tau_{1}\succ\tau_{2})],$
where $\kappa=\sup_{r\in[-R_{max},R_{max}]}\frac{1}{\sigma^{\prime}(r)}$ measures the degree of non-linearity of the sigmoid function. We have applied the mean value theorem, under Assumption 3.2.
As $R_{inf}\in\mathcal{R}$, we have: $P_{R}(\tau_{1}\succ\tau_{2})-P_{\hat{R}_{inf}}(\tau_{1}\succ\tau_{2})\leq u_{P_{R}}(\tau_{1},\tau_{2})$.
Let $d_{pref}$ correspond to the distribution of the preference data, which consists of rollouts from exploratory policies within the learned environment model: $d_{pref}=d^{\pi_{1}}_{\hat{T}_{inf}}/2+d^{\pi_{2}}_{\hat{T}_{inf}}/2$. Recall that the near-optimal policy set $\Pi_{offline}$ includes policy $\pi^{*}_{offline}$ (Lemma A.9) and that $\pi_{1},\pi_{2}$ are the two more exploratory policies within this set:
$\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}_{offline}}_{\hat{T}},\tau_{2}\sim d_{pref}}[u_{P_{R}}(\tau_{1},\tau_{2})]\leq\max_{\pi_{1},\pi_{2}\in\Pi_{offline}}\mathbb{E}_{\tau_{1}\sim d^{\pi_{1}}_{\hat{T},\tau_{2}\sim d^{\pi_{2}}_{\hat{T}}}}[u_{P_{R}}(\tau_{1},\tau_{2})].$
@@
$\mathbb{E}_{(\tau_{1},\tau_{2})\sim\mathcal{D}_{pref}}[\|P_{R}(\tau_{1}\succ\tau_{2})-P_{\tilde{R}}(\tau_{1}\succ\tau_{2})\|_{1}^{2}]\leq\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}{N_{p}},$
@@
$\mathbb{E}_{(\tau_{1},\tau_{2})\sim\mathcal{D}_{pref}}[u_{P_{R}}(\tau_{1},\tau_{2})]\leq 2\sqrt{\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{g}}(1/N_{p})/\delta)}{N_{p}}}.$
@@
$V^{\pi^{*}_{offline}}_{\hat{T}_{inf},R}-V^{\pi^{*}_{offline}}_{\hat{T}_{inf},\hat{R}_{inf}}$ $\leq 2\kappa\sqrt{\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{g}}(1/N_{p})/\delta)}{N_{p}}}.$
@@
$V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{inf},R}\leq HR_{max}C_{T}(\mathcal{F}_{T},\pi^{*})\sqrt{\frac{c_{T}\log(H\mathcal{N}_{\mathcal{F}_{T}}(1/N_{o})/\delta)}{N_{o}}}.$
@@
$V^{\pi^{*}_{offline}}_{\hat{T}_{inf},R}-V^{\pi^{*}_{offline}}_{\hat{T}_{inf},\hat{R}_{inf}}\leq 2\kappa\sqrt{\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}{N_{p}}}.$
@@
Appendix B Implementation Details
@@
For all baselines, transition and reward models were implemented as linear classifiers (for the Star MDP) or as two-layer perceptions with ReLU activation and hidden layer dimension 32 (Gridworld and Sepsis environments). Training was carried out for two or one epochs for the transition and reward models respectively, with the Adam optimizer [Kingma and Ba, 2014] and a learning rate of $10^{-3}$.
@@
1:Observational trajectories dataset $\mathcal{D}_{offline}$. Hyperparameters $\lambda_{T},\lambda_{R}$.
@@
3:Train an ensemble $\mathcal{T}$ of transition models via bootstrapping on the observational data $\mathcal{D}_{offline}$:
$\hat{T}(\cdot|s,a)=\frac{1}{|\mathcal{T}|}\sum_{\tilde{T}\in\mathcal{T}}\tilde{T}(\cdot|s,a);\quad u_{T}(s,a)=\max_{T_{1},T_{2}\in\mathcal{T}}|T_{1}(\cdot|s,a)-T_{2}(\cdot|s,a)|\cdot R_{max}$
4:$\mathcal{D}_{pref}\leftarrow\emptyset$.
@@
$\Pi_{offline}=\{\pi\>|\>\pi=\text{argmax}_{\pi\in\Pi}\mathbb{E}_{(s,a)\sim d^{\pi}_{\hat{T}}}\big{[}\tilde{R}(s,a)-\lambda_{T}u_{T}(s,a)\big{]}\>\forall\tilde{R}\in\mathcal{R}\}$
7:     Identify exploratory policies: $\pi_{1},\pi_{2}=\text{argmax}_{\pi_{1},\pi_{2}\in\Pi_{offline}}\mathbb{E}_{\tau_{1}\sim d^{\pi_{1}}_{\hat{T}},\tau_{2}\sim d^{\pi_{2}}_{\hat{T}}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]$
@@
10:     $\mathcal{D}_{pref}\leftarrow\mathcal{D}_{pref}\cup\{(\tau_{1},\tau_{2},o)\}$.
11:     Train an ensemble $\mathcal{R}$ of reward models via bootstrapping of the preference data $\mathcal{D}_{pref}$:
$\hat{R}(s,a)=\frac{1}{|\mathcal{R}|}\sum_{\tilde{R}\in\mathcal{R}}\tilde{R}(s,a);\quad u_{R}(s,a)=\max_{R_{1},R_{2}\in\mathcal{R}}|R_{1}(\cdot|s,a)-R_{2}(\cdot|s,a)|$
@@
We build $\Pi_{offline}$ by maintaining a policy model for all $\tilde{R}\in\mathcal{R}$, i.e., each element of the reward ensemble. Policy models are optimized to maximize returns under the transition model $\hat{T}$ and the reward function $\tilde{R}-\lambda_{T}u_{T}$ (Sim-OPRL) or $\tilde{R}+\lambda_{T}u_{T}$ (PbOP). Next, the most exploratory policies are identified by generating 10 rollouts of each of the candidate policies within the learned (SimOPRL) or true (PbOP) model. The trajectories $(\tau_{1},\tau_{2})$ maximizing the preference uncertainty function $u_{P_{R}}(\tau_{1},\tau_{2})$ are used for preference feedback. In PbOP, the trajectories are then added to the trajectories buffer and the transition model is retrained for 20 (Star MDP, Gridworld) or 200 steps (Sepsis).
@@
Policy optimization stages, both in estimating optimal policy sets in Sim-OPRL and PbOP and in outputting final policies, are carried out exactly through linear programming for the Star MDP and Gridworld using cvxopt [Diamond and Boyd, 2016], based on code from Lindner et al. [2021], and using Proximal Policy Optimization [Schulman et al., 2017] implemented in stable-baselines3 [Raffin et al., 2021] for the Sepsis environment. In the latter case, after every preference collection episode, reward and policy models were trained from the checkpoint of the previous iteration, for only 20 steps to minimize computation.
@@
Our ablation study for Figure 2(b) is conducted as follows. For Sim-OPRL without pessimism in the output policy, we output the policy that maximizes the value function under the MLE estimate of the transition and reward function, $\hat{T}$ and $\hat{R}$, after preference acquisition. For Sim-OPRL without pessimism in the simulated rollouts, we estimate the optimal policy set $\Pi_{offline}$ in the MLE estimate of the transition model instead of its pessimistic counterpart. Finally, for Sim-OPRL without optimism in the simulated rollouts, we generate rollouts from any two policies in $\Pi_{offline}$ instead of the most explorative ones.
Appendix C Environment Details

### v2

@@ A.5.2 Analysis of the reward term
Let $\Pi_{\textrm{offline}}$ denote the following set of near-optimal pessimistic policies, under the pessimitic transition model $\hat{T}_{\textrm{inf}}$ and the reward confidence set $\mathcal{R}$:
$\Pi_{\textrm{offline}}=\{\pi\>|\>\pi=\text{argmax}_{\pi\in\Pi}\mathbb{E}_{\tau\sim d^{\pi}_{\hat{T}_{\textrm{inf}}}}\big{[}\tilde{R}(\tau)\big{]}\>\forall\tilde{R}\in\mathcal{R}\}$
Under event $\mathcal{E}_{R}$, we have $\pi^{*}_{\textrm{offline}}\in\Pi_{\textrm{offline}}$.
@@
Recall the definition of $\pi^{*}_{\textrm{offline}}$: $\pi^{*}_{\textrm{offline}}=\text{argmax}_{\pi\in\Pi}V_{\hat{T}_{\textrm{inf}},R}^{\pi}$. Note that there is no need to consider the preference baseline term in $V^{\pi}$ when building $\Pi_{\textrm{offline}}$ since it is independent of the policy. Under event $\mathcal{E}_{R}$, we have $R\in\mathcal{R}$. Thus, $\pi^{*}_{\textrm{offline}}\in\Pi_{\textrm{offline}}$. ∎
@@
$V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},R}-V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}}\leq 2\kappa\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)}$
@@
$V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},R}$ $-V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}}$
$=(V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},R}-V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}})+(V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}}-V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})$
$=\mathbb{E}_{\tau\sim d^{{\pi}^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}}}}[R(\tau)]-\mathbb{E}_{\tau\sim d_{\textrm{pref}}}[{R}(\tau)]-\mathbb{E}_{\tau\sim d^{{\pi}^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}}}}[\hat{R}_{\textrm{inf}}(\tau)]+\mathbb{E}_{\tau\sim d_{\textrm{pref}}}[\hat{R}_{\textrm{inf}}(\tau)]$
$=\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}}},\tau_{2}\sim d_{\textrm{pref}}}[R(\tau_{1})-R(\tau_{2})]-\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}}},\tau_{2}\sim d_{\textrm{pref}}}[\hat{R}_{\textrm{inf}}(\tau_{1})-\hat{R}_{\textrm{inf}}(\tau_{2})]$
$\leq\kappa\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}}},\tau_{2}\sim d_{\textrm{pref}}}[P_{R}(\tau_{1}\succ\tau_{2})-P_{\hat{R}_{\textrm{inf}}}(\tau_{1}\succ\tau_{2})],$
where $\kappa=\sup_{r\in[-R_{\textrm{max}},R_{\textrm{max}}]}\frac{1}{\sigma^{\prime}(r)}$ measures the degree of non-linearity of the sigmoid function. We have applied the mean value theorem, under Assumption 3.2.
As $R_{\textrm{inf}}\in\mathcal{R}$, we have: $P_{R}(\tau_{1}\succ\tau_{2})-P_{\hat{R}_{\textrm{inf}}}(\tau_{1}\succ\tau_{2})\leq u_{P_{R}}(\tau_{1},\tau_{2})$.
Let $d_{\textrm{pref}}$ correspond to the distribution of the preference data, which consists of rollouts from exploratory policies within the learned environment model: $d_{\textrm{pref}}=d^{\pi_{1}}_{\hat{T}_{\textrm{inf}}}/2+d^{\pi_{2}}_{\hat{T}_{\textrm{inf}}}/2$. Recall that the near-optimal policy set $\Pi_{\textrm{offline}}$ includes policy $\pi^{*}_{\textrm{offline}}$ (Lemma A.9) and that $\pi_{1},\pi_{2}$ are the two more exploratory policies within this set:
$\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}_{\textrm{offline}}}_{\hat{T}},\tau_{2}\sim d_{\textrm{pref}}}[u_{P_{R}}(\tau_{1},\tau_{2})]\leq\max_{\pi_{1},\pi_{2}\in\Pi_{\textrm{offline}}}\mathbb{E}_{\tau_{1}\sim d^{\pi_{1}}_{\hat{T},\tau_{2}\sim d^{\pi_{2}}_{\hat{T}}}}[u_{P_{R}}(\tau_{1},\tau_{2})].$
@@
$\mathbb{E}_{(\tau_{1},\tau_{2})\sim\mathcal{D}_{\textrm{pref}}}[\|P_{R}(\tau_{1}\succ\tau_{2})-P_{\tilde{R}}(\tau_{1}\succ\tau_{2})\|_{1}^{2}]\leq\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right),$
@@
$\mathbb{E}_{(\tau_{1},\tau_{2})\sim\mathcal{D}_{\textrm{pref}}}[u_{P_{R}}(\tau_{1},\tau_{2})]\leq 2\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)}.$
@@
$V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},R}-V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}}$ $\leq 2\kappa\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)}.$
@@
$V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},R}\leq HR_{\textrm{max}}C_{T}(\mathcal{F}_{T},\pi^{*})\sqrt{\frac{c_{T}}{N_{o}}\log\left(\frac{H}{\delta}\mathcal{N}_{\mathcal{F}_{T}}\left(\frac{1}{N_{o}}\right)\right)}.$
@@
$V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},R}-V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}}\leq 2\kappa\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)}.$
@@
Appendix B Additional Related Work
Table 3: Comparison of broader related work on preference elicitation and preference-based RL.
Framework Offline Policy-based Sampling Robustness Guarantees Practical Implementation
Preference-based Deep RL
   D-REX [Brown et al., 2020]
   PEBBLE [Lee et al., 2021]
   MRN [Liu et al., 2022b]
   SURF [Park et al., 2022]
   PT [Kim et al., 2023]
   IPL [Hejna and Sadigh, 2024]
Preference Elicitation
   PbOP [Chen et al., 2022]
   REGIME [Zhan et al., 2023b]
   FREEHAND [Zhan et al., 2023a]
   OPRL [Shin et al., 2022]
   MoP-RL [Liu et al., 2023]
   Max. Regret [Wilde et al., 2020]
   Biyik and Sadigh [2018]
   Bıyık et al. [2019]
   Sim-OPRL (Ours)
In this section, we position this work against additional related literature. We summarize our analysis in Table 3.
First, we note that significant prior work on preference-based reinforcement learning is centered on improving preference modeling or policy optimization for a fixed (uncertainty-based or uniform) preference elicitation strategy (all methods in the top half of Table 3). This focus is orthogonal to our work, and could certainly be combined with our preference elicitation method for optimal performance.
Next, comparing with other preference elicitation methods in Table 3 (bottom half), we find that almost all works consider a setting with online environment interaction, which makes them incompatible with our problem setting. Online environment exploration eliminates the need for pessimism against transition dynamics [Levine et al., 2020], whereas we show that it is essential for both theoretical and empirical performance in the offline setting.
The uncertainty-based preference elicitation strategy adopted in most practical PbRL algorithms is to minimize uncertainty about preferences, or to maximize information gain about the reward function. From prior theoretical and empirical work [Chen et al., 2022, Lindner et al., 2021], we know that this strategy is often suboptimal compared to actively reducing the set of candidate optimal policies, and both our theoretical and experimental results confirm this.
Finally, note that ours is the first work to provide both a theoretical and empirical analysis of the sample complexity of preference elicitation in the fully offline setting.
Beyond Offline Reinforcement Learning.
With our community’s recent surge in interest in preference-based reinforcement learning for its application to language model alignment [Stiennon et al., 2020, Ouyang et al., 2022], we foresee many possible applications and extensions of our work beyond offline RL. Sample efficiency in preference collection is of paramount importance in Reinforcement Learning from Human Feedback (RLHF): recent works consider active learning pipelines based on uncertainty sampling to improve efficiency [Chen et al., 2024, Melo et al., 2024], similar to OPRL. The insights from our work suggest that a model-based elicitation strategy like Sim-OPRL, designed to simulate responses from near-optimal policies, might perform even better. We believe this line of future work could prove critical to the language model alignment community.
Appendix C Implementation Details
@@
For all baselines, transition and reward models were implemented as linear classifiers (for the Star MDP), as two-layer perceptions with ReLU activation and hidden layer dimension 32 (Gridworld, Sepsis, MiniGrid environments), or as 5-layer MLPs with ReLU activations and hidden sizes $[512,256,128,64,32]$ for the HalfCheetah environments. Training was carried out for two or one epochs for the transition and reward models respectively, with the Adam optimizer [Kingma and Ba, 2014] and a learning rate of $10^{-3}$.
@@
1:Observational trajectories dataset $\mathcal{D}_{\textrm{offline}}$. Hyperparameters $\lambda_{T},\lambda_{R}$.
@@
3:Train an ensemble $\mathcal{T}$ of transition models via bootstrapping on the observational data $\mathcal{D}_{\textrm{offline}}$:
$\hat{T}(\cdot|s,a)=\frac{1}{|\mathcal{T}|}\sum_{\tilde{T}\in\mathcal{T}}\tilde{T}(\cdot|s,a);\quad u_{T}(s,a)=\max_{T_{1},T_{2}\in\mathcal{T}}|T_{1}(\cdot|s,a)-T_{2}(\cdot|s,a)|_{1}\cdot R_{\textrm{max}}$
4:$\mathcal{D}_{\textrm{pref}}\leftarrow\emptyset$.
@@
$\Pi_{\textrm{offline}}=\{\pi\>|\>\pi=\text{argmax}_{\pi\in\Pi}\mathbb{E}_{(s,a)\sim d^{\pi}_{\hat{T}}}\big{[}\tilde{R}(s,a)-\lambda_{T}u_{T}(s,a)\big{]}\>\forall\tilde{R}\in\mathcal{R}\}$
7:     Identify exploratory policies: $\pi_{1},\pi_{2}=\text{argmax}_{\pi_{1},\pi_{2}\in\Pi_{\textrm{offline}}}\mathbb{E}_{\tau_{1}\sim d^{\pi_{1}}_{\hat{T}},\tau_{2}\sim d^{\pi_{2}}_{\hat{T}}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]$
@@
10:     $\mathcal{D}_{\textrm{pref}}\leftarrow\mathcal{D}_{\textrm{pref}}\cup\{(\tau_{1},\tau_{2},o)\}$.
11:     Train an ensemble $\mathcal{R}$ of reward models via bootstrapping of the preference data $\mathcal{D}_{\textrm{pref}}$:
$\hat{R}(s,a)=\frac{1}{|\mathcal{R}|}\sum_{\tilde{R}\in\mathcal{R}}\tilde{R}(s,a);\quad u_{R}(s,a)=\max_{R_{1},R_{2}\in\mathcal{R}}|R_{1}(s,a)-R_{2}(s,a)|_{1}$
@@
We build $\Pi_{\textrm{offline}}$ by maintaining a policy model for all $\tilde{R}\in\mathcal{R}$, i.e., each element of the reward ensemble. Policy models are optimized to maximize returns under the transition model $\hat{T}$ and the reward function $\tilde{R}-\lambda_{T}u_{T}$ (Sim-OPRL) or $\tilde{R}+\lambda_{T}u_{T}$ (PbOP). Next, the most exploratory policies are identified by generating 10 rollouts of each of the candidate policies within the learned (SimOPRL) or true (PbOP) model. The trajectories $(\tau_{1},\tau_{2})$ maximizing the preference uncertainty function $u_{P_{R}}(\tau_{1},\tau_{2})$ are used for preference feedback. In PbOP, the trajectories are then added to the trajectories buffer and the transition model is retrained for 20 (Star MDP, Gridworld) or 200 steps (Sepsis).
@@
Policy optimization stages, both in estimating optimal policy sets in Sim-OPRL and PbOP and in outputting final policies, are carried out exactly through linear programming for the Star MDP and Gridworld using cvxopt [Diamond and Boyd, 2016], based on code from Lindner et al. [2021], using Proximal Policy Optimization [Schulman et al., 2017] implemented in stable-baselines3 [Raffin et al., 2021] for the Sepsis and MiniGrid environments, and Soft Actor-Critic [Haarnoja et al., 2018] for HalfCheetah. In the latter case, after every preference collection episode, reward and policy models were trained from the checkpoint of the previous iteration, for only 20 steps to minimize computation.
@@
Our ablation study for Figure 2 is conducted as follows. For Sim-OPRL without pessimism in the output policy, we output the policy that maximizes the value function under the MLE estimate of the transition and reward function, $\hat{T}$ and $\hat{R}$, after preference acquisition. For Sim-OPRL without pessimism in the simulated rollouts, we estimate the optimal policy set $\Pi_{\textrm{offline}}$ in the MLE estimate of the transition model instead of its pessimistic counterpart. Finally, for Sim-OPRL without optimism in the simulated rollouts, we generate rollouts from any two policies in $\Pi_{\textrm{offline}}$ instead of the most explorative ones.
Appendix D Environment Details